OpenStudy (anonymous):

Solve the equation: x/6+7=20

OpenStudy (jojo4eva):

add 7 to both sides and you get x/6 =27 multiply both sides by 6 and x=162

OpenStudy (the_fizicx99):

x/6 + 7 = 20 -7 -7 x/6 = 13 *6 *6 x = 78 78/6 + 7 = 20 13 + 7 = 20 20 = 20

OpenStudy (jojo4eva):

he is right i added 7 not subtract

OpenStudy (barrelracer011):

First you have to subtract 7 and that would be 13 so your new problem would be x/6=13 then you have to mulitply 6 and 13 and get 78. so x= 78

OpenStudy (jojo4eva):

i made a mistake im sorry

OpenStudy (jojo4eva):

x=78